OpenStudy (anonymous):

The volume of an open box with a square base is 4,000 cu.cm.. Find the dimensions of the box if the material used to make the box is minimun. (Solution musy be with Differential Calculus)

OpenStudy (anonymous):

Recall that the perimeter of a cube : A = l^2 + 4lh but because it is an open box than we will say l = side length V=l^2h so 4000 cm ^3 =l^2h so solve for l \[l = \sqrt{\frac{ 4000 }{ h }}\] Substitute that back into the Area formula. \[A =(\sqrt{\frac{ 4000 }{ h }})^{2} h\] simplify this and take the derivate. Make it = to zero and solve for h. The solve backwards...I think..

OpenStudy (anonymous):

In the last formula, is it "squared"? or "cube"?

OpenStudy (anonymous):

The volume formula is length squared times height.